By contour techniques"\smallint" 1/(2+cos"\theta") limit 0 to 2"\pi" .


1
Expert's answer
2021-09-17T03:44:49-0400

"{\displaystyle\int}\dfrac{1}{\cos\left({\theta}\right)+2}\,\mathrm{d}{\theta}"


"={\displaystyle\int}\dfrac{\sec^2\left(\frac{{\theta}}{2}\right)}{\tan^2\left(\frac{{\theta}}{2}\right)+3}\,\mathrm{d}{\theta}" ...simplify using trigonometric

Substitute 

"u=\dfrac{\tan\left(\frac{{\theta}}{2}\right)}{\sqrt{3}}", "\dfrac{\mathrm{d}u}{\mathrm{d}{\theta}} = \dfrac{\sec^2\left(\frac{{\theta}}{2}\right)}{2\sqrt{3}}"


"\mathrm{d}{\theta}=\dfrac{2\sqrt{3}}{\sec^2\left(\frac{{\theta}}{2}\right)}\,\mathrm{d}u"


"={\displaystyle\int}\dfrac{2\sqrt{3}}{3u^2+3}\,\mathrm{d}u"


"={{\dfrac{2}{\sqrt{3}}}}{\displaystyle\int}\dfrac{1}{u^2+1}\,\mathrm{d}u"


"=\dfrac{2\arctan\left(u\right)}{\sqrt{3}}" substitution u


"=[\dfrac{2\arctan\left(\frac{\tan\left(\frac{{\theta}}{2}\right)}{\sqrt{3}}\right)}{\sqrt{3}}]_0^{2\pi}"


"=\dfrac{2{\pi}}{\sqrt{3}}"